# Universal Ahlfors–David regularity of Steiner trees

### v1

@@ 2.5 Length formula for a planar tree
Let $v_{1},\ldots,v_{n+s}\in\mathbb{C}$ be the vertices of $\mathcal{S}t$ in the complex plane (they include the terminal points $v_{k}=p_{k}$ for $k=1,\dots,n$ and $s\leq n-2$ branching points) and let $E$ be the set of edges of $\mathcal{S}t$. Let $u(z):=\frac{z}{\lvert z\rvert}=e^{i\arg z}$ be the unit complex number representing the direction of the complex number $z$ so that $\lvert z\rvert=z\cdot\overline{u(z)}$. One has
@@
$=\sum_{(k,j)\in E}\left(v_{k}\cdot\overline{u(v_{k}-v_{j})}+v_{j}\cdot\overline{u(v_{j}-v_{k})}\right)=\sum_{k=1}^{n+s}v_{k}\cdot\sum_{(k,j)\in E}\overline{u(v_{k}-v_{j})}.$
@@
The stability of solutions can be combined with the Maxwell-type formula as follows. Let $v_{1},\dots,v_{n}$ be points in the plane for which the Steiner tree $\mathcal{S}t_{v}$ with topology $T$ is unique and full. Let $u_{1},\dots,u_{n}$ be the points obtained by shifting $v_{i}$, such that $|u_{j}-v_{j}|<\varepsilon$ for each $j$. By Lemma 3, for sufficiently small $\varepsilon$, the Steiner tree for the new points $\mathcal{S}t_{u}$ preserves the topology from $D(T)$. For an even smaller $\varepsilon$, the tree $\mathcal{S}t_{u}$ has topology $T$, because the possibility of drawing a full tree with a given topology is determined by a finite system of strict inequalities.
@@
The Steiner ratio of a metric space $X$ is defined as the supremum of the ratio between the length of a minimum spanning tree and that of a Steiner minimal tree over finite sets $\mathcal{A}\subset X$ for which MST exists.
@@
It is worth noting that a direct application of the triangle inequality gives that the Steiner ratio of an arbitrary metric space is at most two.
@@ 2.7 Steiner tree on a surface
Moreover, $c_{d}=2+o_{d}(1)$ and $c_{d}\leq c_{3}=7.76\dots$ for every $d$.
@@
Consider arbitrary points $a_{1},\dots,a_{t}$ on the sphere $\mathbb{S}^{d-1}$; let $\mathcal{S}t$ be a Steiner tree for $\{a_{i}\}$. Let $\varepsilon=ct^{-\frac{1}{d-1}}$, where we will choose the constant $c$ later. Let the angle $\varphi$ be such that $\sin\varphi=\varepsilon$. Consider a maximal (with respect to inclusion) set of points $x_{1},\dots,x_{k}\in\mathbb{S}^{d-1}$ for which the spherical caps $B_{\sin\varphi}(x_{i})\cap\mathbb{S}^{d-1}$ are pairwise disjoint. Then the union of the spherical caps $B_{\sin 2\varphi}(x_{i})\cap\mathbb{S}^{d-1}$ covers the sphere $\mathbb{S}^{d-1}$. Connect the points $x_{1},\dots,x_{k}$ with a spanning tree, and then each point $a_{i}$ with the closest of the points $\{x_{j}\}$. Prim’s algorithm adds an edge no longer than $\sin 4\varphi$ at each iteration, so
$\mathcal{H}^{1}(\mathcal{S}t)\leq\mathcal{H}^{1}(MST)+t\sin 2\varphi\leq(k-1)\sin 4\varphi+t\sin 2\varphi<2(2k+t)\varepsilon,$
since $\sin 4\varphi<4\sin\varphi$ and $\sin 2\varphi<2\sin\varphi$. Estimate $k$; a direct volume estimate gives
$k\leq\frac{\mathcal{H}^{d-1}(B_{\sin\varphi}(x_{i})\cap\mathbb{S}^{d-1})}{\mathcal{H}^{d-1}(\mathbb{S}^{d-1})}\leq\frac{\sqrt{2\pi d}}{\sin^{d-1}\varphi},$
@@
$\mathcal{H}^{1}(\mathcal{S}t)\leq 2\left(2\sqrt{2\pi d}\cdot c^{1-d}+c\right)t^{\frac{d-2}{d-1}}.$
Optimizing over $c$, we get that for $c_{min}=(2\sqrt{2\pi d}(d-1))^{1/d}$
$\mathcal{H}^{1}(\mathcal{S}t)\leq 2\left(2\sqrt{2\pi d}(d-1)\right)^{1/d}\frac{d}{d-1}t^{\frac{d-2}{d-1}}=\left(2+o_{d}(1)\right)t^{\frac{d-2}{d-1}}.$
@@ 2.8 Two bounds on length of a set
Let $S$ be a non-empty measurable subset of $\mathbb{R}^{d}$ and $x\in\mathbb{R}^{d}$ be an arbitrary point. Then
$\mathcal{H}^{1}(S)\geq\int_{r=0}^{\infty}\#(B_{r}(x)\cap S)dr.$
@@ 2.9 Cocircular terminal sets
Suppose that a given set $P=\{p_{1},p_{2},\dots p_{n}\}$ lies on a circle $C\subset\mathbb{R}^{2}$ of radius $r$ with a given cyclic order. If at most one of the segments $p_{i}p_{i+1}]$, $i=1,\dots,n$ (here $n+1=1$) has length strictly greater than $r$, then a Steiner tree for $P$ has no Steiner points.
We need the following corollary.
@@
Let $C\subset\mathbb{R}^{2}$ be a circle of radius $r$. Let $\mathcal{A}\subset C$ be an infinite set such that at most one connected component (arc) of $C\setminus\mathcal{A}$ has diameter greater than $r-\delta$ for some $\delta>0$. Then a solution $\mathcal{S}t$ to the Steiner problem for $\mathcal{A}$ cannot be full.
@@
Now apply Lemma 5 with $X=F$ and $Y=(\partial B_{r}(x)\setminus R_{1}\setminus R_{2})\cup[yz]$. Clearly $\mathcal{H}^{1}(Y)$ is maximal when $\mathcal{H}^{1}([yz])$ is minimal. ∎
@@ 3 Example of a non-regular tree
Here we again identify points in the plane with complex numbers. Consider the unit circle $\omega$. Let $\mathcal{S}t_{0}$ be the Steiner tree for the rectangle $P=\mathcal{A}_{0}$ with vertices $\pm e^{\pi i/6}$, $\pm e^{5\pi i/6}$. It is easy to see that $\mathcal{S}t_{0}$ makes angles of $\pi/3$ with $\omega$ at the vertices. We construct a series of full trees $\mathcal{S}t_{j}$ for terminals $\mathcal{A}_{j}$, whose segments are parallel to the same three straight lines. Moreover, the tree $\mathcal{S}t_{j}$ is unique, and any other locally minimal tree is longer by at least $\delta_{j}>0$.
@@
At the $j$-th iteration, we shift the vertices of the original rectangle $P$ in $\mathcal{S}t_{j-1}$ along the circle $\omega$ by an amount $\varepsilon_{j}<\delta_{j-1}^{2}/64$ inside the angle of size $2\pi/3$, obtaining a new rectangle $P^{\prime}_{j}$. As will be seen later, the directions of the shifts differ between even and odd iterations. Let $\mathcal{S}t^{\prime}_{j}$ be the Steiner tree for $\mathcal{A}^{\prime}_{j}:=\mathcal{A}_{j-1}\cup P^{\prime}_{j}\setminus P$. By Lemma 3, the tree $\mathcal{S}t^{\prime}_{j}$ is a unique Steiner set for the terminal set $\mathcal{A}^{\prime}_{j}$ and has the same topology as $\mathcal{S}t_{j-1}$. By the reasoning after Lemma 3, the tree $\mathcal{S}t^{\prime}_{j}$ is parallel to $\mathcal{S}t_{j-1}$, because if we write the expression for the length with the same coefficients, the difference will be real:
$\sum_{u_{i}\in\mathcal{A}_{j}}\bar{c_{i}}u_{i}-\sum_{u_{i}\in\mathcal{A}^{\prime}_{j}}\bar{c_{i}}u_{i}=2e^{\pi i/6}v+2e^{5\pi i/6}\bar{v}\in\mathbb{R},$
where $v$ is the displacement vector of the vertex $e^{\pi i/6}$. A combinatorial structure of the tree $\mathcal{S}t_{4}$ is depicted in Fig. 4. However, the figure is not drawn to scale; as is evident from the rightmost part of Figure 2, even the transition from $\mathcal{S}t_{1}$ to $\mathcal{S}t_{2}$ cannot be depicted accurately.
$\omega$$\mathcal{S}t_{j-1}$ $x\in P$$\frac{\pi}{3}$
@@
Figure 3: Transition from $\mathcal{S}t_{j-1}$ to $\mathcal{S}t_{j}$ in the neighborhood of a vertex $x\in P$
Now, for each $x\in P$, we replace the edge $x^{\prime}_{j}y_{j}]$ in $\mathcal{S}t^{\prime}_{j}$ with a regular tripod with ends $x,y_{j},z_{j}$ and branching point $t_{j}$, where the point $z_{j}\in\omega$ is chosen so that the tripod intersects $x^{\prime}_{j}y_{j}]$ along a segment. Let $\mathcal{S}t_{j}$ be the resulting tree on terminals $\mathcal{A}_{j}$ with topology $T_{j}$.
Note that
$|xz_{j}|<\sqrt{2\varepsilon_{j}}.$
Suppose that a tree $\mathcal{S}t$ with topology not in $D(T_{j})$ is a Steiner tree for $\mathcal{A}_{j}$. Then
$\mathcal{H}^{1}(\mathcal{S}t)<\mathcal{H}^{1}(\mathcal{S}t_{j})<\mathcal{H}^{1}(\mathcal{S}t_{j-1})+4\sqrt{2\varepsilon_{j}}<\mathcal{H}^{1}(\mathcal{S}t_{j-1})+\delta_{j-1}.$
Now, let us forget that $\mathcal{A}_{j}\setminus\mathcal{A}_{j-1}$ are terminals of $\mathcal{S}t$, which will give us a tree with topology not in $D(T_{j-1})$, shorter than $\mathcal{H}^{1}(\mathcal{S}t_{j-1})+\delta_{j-1}$, a contradiction. By construction, the tree $\mathcal{S}t_{j}$ is full, so by Proposition 1 the tree $\mathcal{S}t_{j}$ is the unique solution to the Steiner problem for $\mathcal{A}_{j}$. Thus, we can define $\delta_{j}>0$ [^1] Calculations show that one can take $\delta_{j}=const\cdot\varepsilon_{l}$, but we do not need the explicit form of the parameters..
@@
Figure 4: A tree with the same topology and symmetries as $\mathcal{S}t_{4}$

### v2

@@ 2.5 Length formula for a planar tree
Let $v_{1},\ldots,v_{2n-2}\in\mathbb{C}$ be the vertices of $\mathcal{S}t$ in the complex plane (they include the terminal points $v_{k}=p_{k}$ for $k=1,\dots,n$ and n2n-2 branching points) and let $E$ be the set of edges of $\mathcal{S}t$. Let $u(z):=\frac{z}{\lvert z\rvert}=e^{i\arg z}$ be the unit complex number representing the direction of the complex number $z$ so that $\lvert z\rvert=z\cdot\overline{u(z)}$. One has
@@
$=\sum_{(k,j)\in E}\left(v_{k}\cdot\overline{u(v_{k}-v_{j})}+v_{j}\cdot\overline{u(v_{j}-v_{k})}\right)=\sum_{k=1}^{2n-2}v_{k}\cdot\sum_{(k,j)\in E}\overline{u(v_{k}-v_{j})}.$
@@
Another corollary concerns a cocircular $\mathcal{A}$.
Corollary 5.
Let $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ be a finite subset of a unit circle $C$. Suppose that $\mathcal{S}t$ is a full Steiner tree for $\mathcal{A}$. Then
$\mathcal{H}^{1}(\mathcal{S}t)=\sum_{i=1}^{n}\cos\beta_{i},$
where $\beta_{i}$ is the angle between the segment of $\mathcal{S}t$ adjacent to $a_{i}$ and the radius $oa_{i}]$.
Proof.
By Lemma 4
$\mathcal{H}^{1}(\mathcal{S}t)=\sum_{i=1}^{n}e^{\pm\beta_{i}},$
where a sign depends on which of the two possible rays forming an angle $\beta_{i}$ with the radius $oa_{i}]$ contains the segment of $\mathcal{S}t$. Since $\mathcal{H}^{1}(\mathcal{S}t)$ is a real number, we have
$\mathcal{H}^{1}(\mathcal{S}t)=\Re\sum_{i=1}^{n}e^{\pm\beta_{i}}=\sum_{i=1}^{n}\Re e^{\pm\beta_{i}}=\sum_{i=1}^{n}\cos\beta_{i}.$
Next, the stability of solutions can be combined with the Maxwell-type formula as follows. Let $v_{1},\dots,v_{n}$ be points in the plane for which the Steiner tree $\mathcal{S}t_{v}$ with topology $T$ is unique and full. Let $u_{1},\dots,u_{n}$ be the points obtained by shifting $v_{i}$, such that $|u_{j}-v_{j}|<\varepsilon$ for each $j$. By Lemma 3, for sufficiently small $\varepsilon$, the Steiner tree for the new points $\mathcal{S}t_{u}$ preserves the topology from $D(T)$. For an even smaller $\varepsilon$, the tree $\mathcal{S}t_{u}$ has topology $T$, because the possibility of drawing a full tree with a given topology is determined by a finite system of strict inequalities.
@@
The Steiner ratio of a metric space $X$ is defined as the supremum of the ratio between the length of a minimum spanning tree and that of a Steiner minimal tree over finite sets $\mathcal{A}\subset X$.
@@
It is worth noting that a direct application of the triangle inequality gives that the Steiner ratio of an arbitrary metric space is at most two, see [3].
@@ 2.7 Steiner tree on a surface
Moreover, $c_{d}=2+o_{d}(1)$ and $c_{d}\leq c_{3}=11.78\dots$ for every $d$.
@@
Consider arbitrary points $a_{1},\dots,a_{t}$ on the sphere $\mathbb{S}^{d-1}$; let $\mathcal{S}t$ be a Steiner tree for $\{a_{i}\}$. Let $\varepsilon=ct^{-\frac{1}{d-1}}$, where we will choose the constant $c$ later. Let the angle $\varphi$ be such that $\sin\varphi=\varepsilon$ and put $\delta=2\sin(\varphi/2)$. This gives that a spherical cap $B_{\delta}(x)\cap\mathbb{S}^{d-1}$, where $x\in\mathbb{S}^{d-1}$ has angular radius $\varphi$, i.e. consists of all points $x^{\prime}\in\mathbb{S}^{d-1}$ such that $\angle xox^{\prime}<\varphi$.
Consider a maximal (with respect to inclusion) set of points $x_{1},\dots,x_{k}\in\mathbb{S}^{d-1}$ for which the spherical caps $B_{\delta}(x_{i})\cap\mathbb{S}^{d-1}$ are pairwise disjoint. Then the union of the spherical caps $B_{2\sin\varphi}(x_{i})\cap\mathbb{S}^{d-1}$ covers the sphere $\mathbb{S}^{d-1}$. Connect the points $x_{1},\dots,x_{k}$ with a spanning tree, and then each point $a_{i}$ with the closest of the points $\{x_{j}\}$. Prim’s algorithm adds an edge no longer than $2\sin 2\varphi$ at each iteration, so
$\mathcal{H}^{1}(\mathcal{S}t)\leq\mathcal{H}^{1}(MST)+2t\sin\varphi\leq 2(k-1)\sin 2\varphi+2t\sin\varphi<2(2k+t)\varepsilon,$
since $\sin 2\varphi<2\sin\varphi$. Estimate $k$; a direct volume estimate gives
$k\leq\frac{\mathcal{H}^{d-1}(\mathbb{S}^{d-1})}{\mathcal{H}^{d-1}(B_{\delta}(x_{i})\cap\mathbb{S}^{d-1})}\leq\frac{\sqrt{2\pi d}}{\sin^{d-1}\varphi}=\frac{\sqrt{2\pi d}}{\varepsilon^{d-1}},$
@@
$\mathcal{H}^{1}(\mathcal{S}t)\leq 2\left(2\sqrt{2\pi d}\cdot c^{2-d}+c\right)t^{\frac{d-2}{d-1}}.$
Optimizing over $c$, we get the following upper bound, which is obtained by the substitution of $c_{min}=(2\sqrt{2\pi d}(d-2))^{1/(d-1)}$
$\mathcal{H}^{1}(\mathcal{S}t)\leq 2\left(2\sqrt{2\pi d}(d-2)\right)^{1/(d-1)}\frac{d-1}{d-2}t^{\frac{d-2}{d-1}}=\left(2+o_{d}(1)\right)t^{\frac{d-2}{d-1}}.$
A numerical analysis of small cases finishes the proof. ∎
@@ 2.8 Two bounds on length of a set
Let $S$ be a nonempty measurable subset of $\mathbb{R}^{d}$ and $x\in\mathbb{R}^{d}$ be an arbitrary point. Then
$\mathcal{H}^{1}(S)\geq\int_{r=0}^{\infty}\#(\partial B_{r}(x)\cap S)dr.$
@@ 2.9 Cocircular terminal sets
Suppose that a given set $P=\{p_{1},p_{2},\dots,p_{n}\}$ lies on a circle $C\subset\mathbb{R}^{2}$ of radius $r$ with a given cyclic order. If at most one of the segments $p_{i}p_{i+1}]$, $i=1,\dots,n$ (here $n+1=1$) has length strictly greater than $r$, then a Steiner tree for $P$ has no Steiner points.
We need to adapt the result for the case of infinite $\mathcal{A}$.
@@
Let $C\subset\mathbb{R}^{2}$ be a circle of radius $r$. Let $\mathcal{A}\subset C$ be an infinite compact set such that at most one connected component (arc) of $C\setminus\mathcal{A}$ has diameter greater than $r-\delta$ for some $\delta>0$. Then a solution $\mathcal{S}t$ to the Steiner problem for $\mathcal{A}$ has no Steiner points.
@@
Now apply Lemma 5 with $X=F$ and $Y=(\partial B_{r}(x)\setminus R_{1}\setminus R_{2})\cup[yz]$. Clearly $\mathcal{H}^{1}(Y)$ is maximal when $\mathcal{H}^{1}([yz])$ is minimal. If $|yz|=r$, then $\mathcal{H}^{1}(Y)=4\pi r/3+r$. ∎
@@ 3 Example of a non-regular tree
Here we again identify points in the plane with complex numbers. Consider the unit circle $\omega$. Let $\mathcal{S}t_{0}$ be the Steiner tree for the set $\mathcal{A}_{0}=\{\pm e^{\pi i/6},\pm e^{5\pi i/6}\}$, whose points form rectangle $P$, see the leftmost part of Fig. 2. It is easy to see that $\mathcal{S}t_{0}$ makes angles of $\pi/3$ with $\omega$ at the vertices. We construct a sequence of full trees $\mathcal{S}t_{j}$ for terminals $\mathcal{A}_{j}$, whose edges are parallel to the same three straight lines. Moreover, every tree $\mathcal{S}t_{j}$ is unique Steiner tree for $\mathcal{A}_{j}$, and any other locally minimal tree for $\mathcal{A}_{j}$ is longer than $\mathcal{S}t_{j}$ by at least $\delta_{j}>0$, which will be defined later.
@@
At the $j$-th iteration, we shift the terminals from $\mathcal{A}_{0}$ in $\mathcal{A}_{j-1}$ (by construction $\mathcal{A}_{0}\subset\mathcal{A}_{j-1}$ for every $j$) along the circle $\omega$ by an amount $\varepsilon_{j}<\delta_{j-1}^{2}/128j$ inside the angle of size $2\pi/3$, obtaining a new rectangle with the vertex set $P_{j}$. Formally, the direction can be defined by the decreasing of $(-1)^{j}|p_{y}|$, where $p_{y}$ is the $y$-coordinate of the point $p\in\mathcal{A}_{0}$, so the directions of the shifts differ between even and odd iterations. Let $\mathcal{S}t^{\prime}_{j}$ be the Steiner tree for $\mathcal{A}^{\prime}_{j}:=\mathcal{A}_{j-1}\cup P_{j}\setminus\mathcal{A}_{0}$. By Lemma 3, the tree $\mathcal{S}t^{\prime}_{j}$ is a unique Steiner set for the terminal set $\mathcal{A}^{\prime}_{j}$ and has the same topology as $\mathcal{S}t_{j-1}$.
Analogously to the reasoning from the end of Subsection 2.5, we will show that the tree $\mathcal{S}t^{\prime}_{j}$ is parallel to $\mathcal{S}t_{j-1}$. Indeed, let $\{c_{i}\}$ be the set of coefficients in Maxwell-type formula (1) for $\mathcal{S}t_{j-1}$. Write the difference of (1) for $\mathcal{S}t_{j-1}$ and the same expression with the same coefficients for the terminal set $\mathcal{A}^{\prime}_{j}$
$\Delta_{j}:=\sum_{u\in\mathcal{A}_{j-1}}\bar{c_{u}}u-\sum_{u\in\mathcal{A}^{\prime}_{j}}\bar{c_{u}}u=\sum_{u\in\mathcal{A}_{0}}\bar{c_{u}}u-\sum_{u\in P_{j}}\bar{c_{u}}u.$
In the case of odd $j$ the coefficients $\bar{c_{u}}$ for $e^{\pi i/6}$, $e^{5\pi i/6}$, $e^{7\pi i/6}$ and $e^{11\pi i/6}$ are $e^{-\pi i/3}$, $e^{-2\pi i/3}$, $e^{2\pi i/3}$ and $e^{\pi i/3}$, respectively. In this case
$\Delta_{j}=-2e^{-i\pi/3}v_{j}-2e^{i\pi/3}\bar{v_{j}},$
where $v_{j}$ is the displacement vector of the vertex $e^{\pi i/6}\in\mathcal{A}_{0}$. Clearly $\Delta_{j}$ is real, so the trees $\mathcal{S}t^{\prime}_{j}$ and $\mathcal{S}t_{j-1}$ are segment-wise parallel. The case of even $j$ is completely analogous with
$\Delta_{j}=-2v_{j}-2\bar{v_{j}}.$
Now, for each $x\in\mathcal{A}_{0}$ denote by $x_{j}$ the corresponding shifted terminal in $\mathcal{S}t^{\prime}_{j}$ (so for $x=e^{\pi i/6}$ we have $x_{j}=x+v_{j}$). To finalize the $j$-th step we replace for every $x\in\mathcal{A}_{0}$ the edge $x_{j}y(x)_{j}]$ in $\mathcal{S}t^{\prime}_{j}$ with a regular tripod with ends $x,y(x)_{j},z(x)_{j}$ and branching point $t(x)_{j}$, where the point $z(x)_{j}\in\omega$ is chosen so that the tripod intersects $x_{j}y(x)_{j}]$ along a segment. Let $\mathcal{S}t_{j}$ be the resulting tree on terminals $\mathcal{A}_{j}$ with topology $T_{j}$; so $\mathcal{A}_{j}$ is the union of $\mathcal{A}_{j-1}$ and four points $z(x)_{j}$.
$\omega$$\mathcal{S}t_{j-1}$ $x\in\mathcal{A}_{0}$$\frac{\pi}{3}$
@@
Figure 3: Transition from $\mathcal{S}t_{j-1}$ to $\mathcal{S}t_{j}$ in the neighborhood of vertices $x=e^{\pi i/6},e^{7\pi i/6}$ in the case of even $j$ and $x=e^{5\pi i/6},e^{11\pi i/6}$ in the case of odd $j$
The combinatorial structure of the tree $\mathcal{S}t_{4}$ is depicted in Fig. 4. However, the figure is not drawn to scale; as is evident from the rightmost part of Figure 2, even the transition from $\mathcal{S}t_{1}$ to $\mathcal{S}t_{2}$ cannot be depicted accurately.
The distance from $t(x)_{j}$ to $\omega$ is at most $\varepsilon_{j}$, and so
$|xz(x)_{j}|<2|t(x)_{j}z(x)_{j}|<2\sqrt{2\varepsilon_{j}}.$
Suppose that a tree $\mathcal{S}t$ with topology $R$ not in $D(T_{j})$ is a Steiner tree for $\mathcal{A}_{j}$. Then
$\mathcal{H}^{1}(\mathcal{S}t)\leq\mathcal{H}^{1}(\mathcal{S}t_{j})<\mathcal{H}^{1}(\mathcal{S}t_{j-1})+8\sqrt{2\varepsilon_{j}}<\mathcal{H}^{1}(\mathcal{S}t_{j-1})+\delta_{j-1}.$
Now, let us forget that vertices $P_{j}=\mathcal{A}_{j}\setminus\mathcal{A}_{j-1}$ are terminals of $\mathcal{S}t$: if a vertex $p\in P_{j}$ has degree 1, then we delete $p$ from the topology $R$ (viewed as an abstract graph) and in the case that $p$ has degree two we contract the adjacent edges $\{p,v_{1}\}$, $\{p,v_{2}\}$ into $\{v_{1},v_{2}\}$ and delete the vertex $p$. Let $\mathcal{S}t^{\prime}$ be the unique (by Proposition 1) local minimizer for the reduced topology $R^{\prime}$. Since $\mathcal{H}^{1}(\mathcal{S}t^{\prime})\leq\mathcal{H}^{1}(\mathcal{S}t)<\mathcal{H}^{1}(\mathcal{S}t_{j-1})+\delta_{j-1}$, the definition of $\delta_{j-1}$ implies $R^{\prime}\in D(T_{j-1})$. Thus the only possibility for $R\notin D(T_{j})$ is that there exist point $p\in P_{j}$ that is adjacent to a point $q$ that is neither the corresponding point to $p_{0}\in\mathcal{A}_{0}$ nor a common branching point with $p_{0}$. But in this case $|pq|>|pp_{0}|$, which is impossible in a Steiner tree. This is a contradiction.
By construction, the tree $\mathcal{S}t_{j}$ is full, so by Proposition 1 the tree $\mathcal{S}t_{j}$ is the unique solution to the Steiner problem for $\mathcal{A}_{j}$. Thus, we can define $\delta_{j}>0$ [^1] Calculations show that one can take $\delta_{j}=const\cdot\varepsilon_{j}$, but we do not need the explicit form of the parameters..
@@
Figure 4: A locally minimal tree with the same topology and symmetries as $\mathcal{S}t_{4}$